OpenStudy (perfecta):

A ball having a mass of 500 grams is dropped from a height of 9.00 meters. What is its kinetic energy when it hits the ground?

OpenStudy (adonine):

ok lets see

OpenStudy (adonine):

44.1 joules

OpenStudy (adonine):

M*g*h where m = mass, g = acceleration due to gravity (9.8 m/s^2), h = height Given m = 500g = 0.5 kg, h = 9 meters 0.5*9*9.8 = 44.1 joules